OpenStudy (anonymous):

Together, teammates pedro and ricky got 2679 base hits last season pedro had 281 more hits than ricky. How many did each player have

OpenStudy (anonymous):

Pls help tnx

OpenStudy (anonymous):

Hits of Pedro + Hits of Ricky = 2679 Let Hits of Ricky = X and Hits of Pedro = X+281 (X + 281) + X = 2679 X + 281 + X = 2679 2X = 2679 - 281 2X = 2398 2X/2 = 2398/2 X = 1199 Hits of Ricky = X = 1199 Hits Hits of Pedro = X+281 = 1199 + 281 = 1480 Hits Check: 1199 + 1480 = 2679 2679 = 2679 Hope this helped! :)

OpenStudy (anonymous):

Thanks for the help!